When 0.85mol of zinc nitrate react with 0.254mol of sodium phosphate insoluble zinc phosphate forms. Which is the limiting reactant and which is the reactant in excess?


1
Expert's answer
2021-06-07T03:24:10-0400

Solution:

zinc nitrate = Zn(NO3)2

sodium phosphate = Na3PO4

zinc phosphate = Zn3(PO4)2


Balanced chemical equation:

3Zn(NO3)2(aq) + 2Na3PO4(aq) → Zn3(PO4)2(s) + 6NaNO3(aq)

According to stoichiometry:

3 mol of Zn(NO3)2 reacts with 2 mol of Na3PO4

Thus 0.85 mol of Zn(NO3)2 reacts with:

(0.85 mol Zn(NO3)2) × (2 mol Na3PO4 / 3 mol Zn(NO3)2) = 0.57 mol Na3PO4

However, initially there is 0.254 mol of Na3PO4 (according to the task).

Thus, Na3PO4 acts as limiting reactant and Zn(NO3)2 is excess reactant.


Answer:

Na3PO4 is the limiting reactant.

Zn(NO3)2 is the reactant in excess.